# From Perceptions to Decisions: Wildfire Evacuation Decision Prediction with Behavioral Theory-informed LLMs

### v1

@@ 4.1 Main Results
In this section, we present a comparative analysis between our approach and several widely adopted methods for wildfire evacuation decision prediction. Specifically, we perform experiments on a consistent dataset to showcase the superior performance of our method, and we further validate its generalizability through cross-dataset testing.
In the consistent dataset experiments (see Table 1), we compare the performance of our method, FLARE—which employs three distinct backbones (GPT-o3-mini OpenAI (2025), GPT-4o Hurst et al. (2024), and Claude-3.5 Anthropic (2024)) separately—with four widely adopted prediction methods: Logistic Regression, Random Forest, LLM Inference with GPT-4o, and HCM. These comparisons focus on predicting evacuation decisions on a combined dataset. The combined dataset is curated by merging survey data from multiple wildfire events. Detailed results for the single-event dataset are provided in Appendix C. The results consistently demonstrate that FLARE achieves superior accuracy in evacuation prediction compared to the baseline methods. In contrast, the baseline approaches not only deliver lower overall accuracy but also struggle with balanced detection across various predictions, as evidenced by their F1 scores. Moreover, FLARE exhibits notable adaptability across different state-of-the-art LLMs, consistently enhancing performance when employing various backbones. Notably, when using Claude-3.5 as the backbone, FLARE improves accuracy by 13.2%, Macro F1 by 12.7%, and Weighted F1 by 11.9%. These improvements indicate that advancements in LLM reasoning capabilities Anthropic (2024) could further elevate the performance of FLARE.
In the cross-dataset generalization tests (see Table 2), FLARE consistently outperforms the baseline methods. The experiments were conducted using the Kincade Fire and Marshall Fire datasets in a cross-validation setup, where one dataset served as the training set and the other as the test set. This setup is designed to account for the fact that the wildfire occurred in two different states, as illustrated in Appendix D. Within the same state, evacuation laws and processes are generally similar for such events. By considering distinct states, this setup maximizes the differences between wildfire events, allowing for a more rigorous evaluation of the generalizability of our proposed methods. As shown in Table 2, FLARE achieved superior performance in terms of accuracy, Macro F1, and weighted F1 scores across both configurations. Notably, the baseline methods—particularly Logistic Regression and Random Forest—struggled to accurately classify the “Stay” class. In contrast, FLARE delivered higher performance metrics and maintained a more balanced detection across classes. These results underscore the robustness of FLARE in cross-event scenarios, highlighting its potential for effective knowledge transfer between different wildfire events.
@@ 5 Discussion - Why It Works
The framework’s effectiveness is driven by the meticulous design of each component, enabling LLMs to generate accurate evacuation predictions and perform complex reasoning based on survey data. One critical aspect is the use of CoT reasoning is crucial for the prediction. Research shows that without carefully designed instructions, LLMs struggle to grasp human mind states, underscoring the need for advanced mechanisms to improve real-world understanding Xu et al. (2024a). Moreover, the evacuation decision process demands complex reasoning, and CoT enables LLMs to mimic human problem-solving—effectively addressing this complexity while enhancing accuracy and robustness Wei et al. (2022); Wang et al. (2022); Kojima et al. (2022). Thus, our CoT design integrates both inferred perceptions and external information, establishing a strong foundation for accurate evacuation decision prediction.
Another key component is the design of the RL module, which aligns the reasoning process with human thoughts. Although LLMs can closely approximate human performance in mental state prediction with well-crafted CoT, their accuracy drops in behavior prediction without detailed cues Gu et al. (2024). However, providing extensive detail increases the input sequence length, straining the LLM’s limited attention and memory and disrupting the coherent CoT required for accurate multi-step reasoning Li et al. (2024c, f); Levy et al. (2024); Qian et al. (2024). To address this, our memory mechanism stores and retrieves the most similar error records during reasoning, ensuring the model obtains sufficient context from previous errors without overextending the input sequence. Moreover, the LLM’s self-reflection capability enables it to autonomously correct errors and refine its responses, thereby enhancing the quality of its output rationale without human intervention Renze and Guven (2024); Li et al. (2023, 2024b); Shinn et al. (2024). When a prediction is incorrect, our self-reflection process ensures that high-quality reflective rationales are stored in Memory, making the contextual cues even more valuable.
Furthermore, the LLM’s natural language understanding ability plays a crucial role in effective prediction. As noted earlier, sufficient information and cues lead to accurate behavior prediction Gu et al. (2024). For example, survey responses that include self-reported medical conditions (e.g., “Extremely disabled from stroke”) are difficult for traditional models to incorporate. In contrast, the LLM can both utilize this information and understand its implications, resulting in more accurate predictions Xie (2011); Qin et al. (2024).
@@ 6 Conclusion
In this study, we introduced FLARE, a novel framework that integrates the reasoning capabilities of LLMs with a well-established behavior theory to address the complexities of human mental states in wildfire evacuation decision prediction. By systematically classifying variables and constructing CoT grounded in threat assessments and risk perceptions, our approach captures evacuees’ heterogeneous preferences and interprets essential perceptions. Moreover, we integrate a memory-based RL module that serves as a dynamic repository of previous errors and justifications, guiding the LLM toward spontaneously improved reasoning. Comprehensive experiments on real-world survey data from historical wildfire events demonstrate that FLARE not only outperforms established methods but also maintains robust generalizability across different wildfire events. Overall, FLARE broadens the innovative applications of LLMs.
@@ Appendix B Prompt Design
Figure 2: The prompt used to generate threat assessment
@@
Figure 3: The prompt used to generate risk perception.
@@
Figure 4: The complete CoT used of evacuation decision prediction.
@@ Appendix C More result
The accuracy of heatmaps for risk perceptions and threat assessment predictions reveals key trends in the model’s performance. Overall, the LLM demonstrates moderate accuracy, with better performance in predicting mid-range values (scores 2–4) while struggling with extreme values (scores 1 and 5). For instance, in risk perception prediction, the model performs best when the actual values are within the 2–4 range, with the highest accuracy (80%) observed when the actual risk perception is 5, but the model predicts 3, indicating a systematic underestimation of extreme risk perceptions. Similarly, in threat assessment prediction, the model achieves its highest accuracy (50%) when the actual threat assessment is 1, frequently predicting 2 instead. This pattern suggests that the model is biased toward moderate assessments and struggles to distinguish individuals with extremely high or low-risk perceptions or threat assessments.
@@ D.3 Marshall Fire
The Marshall Fire ignited shortly before 10:30 a.m. on December 30, 2021, in Boulder County, Colorado, from two distinct ignition points.
Figure 8: The wildfire impact area and evacuation area map of Marshall Fire Forrister et al. (2024a)
From these points, the fire rapidly spread into several suburban areas across Boulder, Jefferson, and Adams Counties, affecting counties such as Louisville, Superior, Broomfield, Lafayette, Arvada, and Westminster. The preceding unusually humid spring, followed by a warm, dry summer and fall, produced dry fuel conditions that, combined with strong dry winds, accelerated the fire’s spread. Recognized as the most destructive wildfire in Colorado’s history, the Marshall Fire burned over 6,200 acres, destroyed 1,084 homes, damaged 149 additional homes, and resulted in two fatalities. In Boulder County alone, the estimated residential damage exceeded \$513 million. On the day of the fire, more than 30,000 residents in Boulder, Jefferson, and Adams Counties were evacuated. Figure 8 illustrates the affected evacuation and pre-evacuation zones for these communities Forrister et al. (2024a).

### v2

@@ 4 Experiment
Additionally, beyond evaluating predictive accuracy, it is also important to assess the reasoning processes generated by the LLM. To this end, we compare them with open-ended questions interviewees answered in the survey (answered by 206 out of 604 interviewees), which could be treated as partial ground truth for reasoning processes. The detailed comparison can be found in Appendix G.
@@ 4.1 Main Results
We evaluate our proposed method, FLARE, against several widely used approaches for wildfire evacuation decision prediction. We first conduct experiments on a consistent dataset to assess overall performance, followed by cross-dataset evaluations to test generalizability.
In the consistent dataset experiments (see Table 1), we compare the performance of our method, FLARE—which employs three distinct backbones (GPT-o3-mini OpenAI (2025), GPT-4o Hurst et al. (2024), and Claude-3.5 Anthropic (2024)) separately—with four widely adopted prediction methods: Logistic Regression, Random Forest, LLM Inference with GPT-4o, and HCM. The dataset was constructed by merging survey responses from multiple wildfire events. Results on individual datasets are provided in Appendix C.
The results consistently demonstrate that FLARE achieves superior accuracy in evacuation prediction compared to the baseline methods. In contrast, the baseline approaches not only deliver lower overall accuracy but also struggle with balanced detection across various predictions, as evidenced by their F1 scores. Moreover, FLARE exhibits notable adaptability across different state-of-the-art LLMs, consistently enhancing performance when employing various backbones. Notably, when using Claude-3.5 as the backbone, FLARE improves accuracy by 13.2%, Macro F1 by 12.7%, and Weighted F1 by 11.9%. These improvements indicate that advancements in LLM reasoning capabilities Anthropic (2024) could further elevate the performance of FLARE.
In the cross-dataset generalization experiments (see Table 2), we use the Kincade Fire and Marshall Fire datasets in a cross-validation setup, where one dataset served as the training set and the other as the test set. This setup is designed to account for the fact that the wildfire occurred in two different states, as illustrated in Appendix D. Within the same state, evacuation laws and processes are generally similar for such events. By considering distinct states, this setup maximizes the differences between wildfire events, allowing for a more rigorous evaluation of the generalizability of our proposed methods. As shown in Table 2, FLARE achieved superior performance in terms of accuracy, Macro F1, and weighted F1 scores across both configurations. Notably, the baseline methods—particularly Logistic Regression and Random Forest—struggled to accurately classify the “Stay” class. In contrast, FLARE delivered higher performance metrics and maintained a more balanced detection across classes. These results underscore the robustness of FLARE in cross-event scenarios, highlighting its potential for effective knowledge transfer between different wildfire events.
@@ 5 Discussion - Why It Works
The framework’s effectiveness is driven by the meticulous design of each component, enabling LLMs to generate accurate evacuation predictions through complex behavioral patterns extracted from survey data. A key innovation in our framework is the integration of a reasoning pattern classifier with behaviorally grounded Chain-of-Thought (CoT) prompting. Prior studies show that LLMs struggle to infer human mental states without structured guidance Xu et al. (2024a). To address this, we first use a classifier to identify each individual’s most likely reasoning pattern based on behavioral constructs such as threat assessment and risk perception. The selected pattern then guides the construction of a customized CoT template. This design grounds LLM reasoning in psychological theory while adapting it to individual decision logic. The resulting CoT incorporates classifier-derived perceptions and external context, enabling the LLM to emulate complex human reasoning Wei et al. (2022); Wang et al. (2022); Kojima et al. (2022) and produce more accurate, interpretable evacuation predictions.
Another key component of our framework is the memory-based reinforcement learning (RL) mechanism, which integrates error correction and self-reflection to align LLM reasoning with human behavior. While LLMs can approximate mental state inference with structured CoT, their behavior prediction often degrades without detailed contextual cues Gu et al. (2024). Yet providing excessive detail can overwhelm the model’s attention span and disrupt coherent reasoning Li et al. (2024b); Levy et al. (2024); Qian et al. (2024). To mitigate this, we store past inference errors along with their CoT rationales and self-generated reflections. During prediction, the model retrieves similar error cases from memory to provide targeted, context-rich cues without inflating input length. This reflective process allows the LLM to iteratively refine its reasoning, enhancing both prediction accuracy and interpretability Renze and Guven (2024); Li et al. (2023a, 2024a); Shinn et al. (2024).
Additionally, our framework performs perception identification by extracting individual threat and risk perceptions through PADM-based variable selection Sun et al. (2024c). This process provides structured inputs for reasoning and grounds the CoT in established behavioral theory, ensuring that the model’s inferences reflect psychologically meaningful mental states rather than generic patterns. This alignment further strengthens the connection between internal perceptions and observable evacuation behavior.
@@ 6 Conclusion
In this study, we introduced FLARE, a novel framework that integrates the reasoning capabilities of LLMs with a well-established behavioral theory to predict the complexities of human mental states and evacuation decisions in wildfire. By systematically classifying variables and building CoT grounded in threat assessments and risk perceptions, our approach captures evacuees’ heterogeneous preferences and interprets essential perceptions. Moreover, we integrate a memory-based RL module that serves as a dynamic repository of previous errors and justifications, guiding the LLM toward spontaneously improved reasoning. Comprehensive experiments on real-world survey data from historical wildfire events demonstrate that FLARE not only outperforms established methods but also maintains robust generalizability across different wildfire events. Overall, FLARE introduces a novel framework to effectively integrate behavioral theory to inform and improve LLMs expressive reasoning capabilities. It enables a rigorous, theory-constrained alignment between LLM CoTs and human reasoning process, broadening the innovative use of LLMs in mimicking and predicting human behaviors.
@@
Acknowledgements
This research was supported by the National Science Foundation (Award #2442712, #2332145, and #2338959), and by the Rural Equitable and Accessible Transportation (REAT) Center funded by the U.S. Department of Transportation. This research was also supported by the Natural Hazards Center Quick Response Research Program, which is based on work supported by the National Science Foundation (Award #1635593). Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of NSF, USDOT, or the Natural Hazards Center.
@@ Appendix B Prompt Design
Table 5: The prompt used to generate threat assessment
@@
Table 6: The prompt used to generate risk perception.
@@
Table 7: The complete CoT used of evacuation decision prediction.
@@ Appendix C More result
C.1 Accuracy heatmaps for risk and threat perception predictions
The accuracy heatmaps for risk perceptions and threat assessment predictions reveals key trends in the model’s performance. Overall, the LLM demonstrates moderate accuracy, with better performance in predicting mid-range values (scores 2–4) while struggling with extreme values (scores 1 and 5). For instance, in risk perception prediction, the model performs best when the actual values are within the 2–4 range, with the highest accuracy (80%) observed when the actual risk perception is 5, but the model predicts 3, indicating a systematic underestimation of extreme risk perceptions. Similarly, in threat assessment prediction, the model achieves its highest accuracy (50%) when the actual threat assessment is 1, frequently predicting 2 instead. This pattern suggests that the model is biased toward moderate assessments and struggles to distinguish individuals with extremely high or low-risk perceptions or threat assessments.
@@ C.1 Accuracy heatmaps for risk and threat perception predictions
C.2 Further exploration in threshold settings
To further explore justify the necessity of selecting representative variable subsets for each perception, we conducted an experiment using all survey variables without selection. The results in Table 9 demonstrate a notable performance degradation compared to the original variable selection strategy. This drop is due to the transformer’s limited capacity to retain and reason over long, dense input sequences—when overwhelmed with less relevant information, its attention becomes diluted and reasoning less coherent.
C.3 Detailed Comparison with Supervised Fine-tuning methods
As shown in Table 11, the SFT model achieved moderate performance in within-dataset settings (e.g., 0.7667 accuracy on Kincade and 0.724 on Marshall). However, it struggled significantly in cross-dataset generalization, especially when training on Kincade and testing on Marshall. In contrast, our framework demonstrates consistent and superior performance across all configurations. These results suggest that supervised fine-tuning alone cannot effectively model the reasoning process behind evacuation behavior, whereas our structured reasoning framework grounded in behavioral theory provides better robustness and transferability.
Train/Test Set Accuracy Macro F1 Weighted F1
Kincade / Kincade 0.767 0.652 0.775
Marshall / Marshall 0.724 0.725 0.724
Marshall / Kincade 0.733 0.544 0.725
Kincade / Marshall 0.6704 0.6703 0.6698
Combined / Combined 0.738 0.557 0.752
Table 9: Performance using all variables (no selection)
Train/Test Set Accuracy Macro F1 Weighted F1
Kincade / Kincade 0.8025 0.7504 0.8222
Marshall / Marshall 0.8222 0.8219 0.8231
Marshall / Kincade 0.7284 0.6680 0.7569
Kincade / Marshall 0.7556 0.7556 0.7556
Combined / Combined 0.8158 0.8021 0.8240
Table 10: Performance under original setting (thresholded variable selection)
Method Train/Test Set Accuracy Macro F1 Weighted F1
SFT Kincade / Kincade 0.7667 0.652 0.775
SFT Marshall / Marshall 0.724 0.7245 0.724
SFT Marshall / Kincade 0.733 0.544 0.725
SFT Kincade / Marshall 0.5714 0.5453 0.5343
SFT Combined / Combined 0.738 0.557 0.752
FLARE Kincade / Kincade 0.803 0.750 0.822
FLARE Marshall / Marshall 0.822 0.821 0.823
FLARE Marshall / Kincade 0.728 0.668 0.757
FLARE Kincade / Marshall 0.756 0.7556 0.756
FLARE Combined / Combined 0.816 0.802 0.824
Table 11: Comparison of Supervised Fine-Tuning (SFT) and Our Method Across Datasets
@@ D.3 Marshall Fire
The Marshall Fire ignited shortly before 10:30 a.m. on December 30, 2021, in Boulder County, Colorado, from two ignition points.
Figure 5: The wildfire impact and evacuation zones of the Marshall Fire Forrister et al. (2024a).
The fire quickly spread into suburban areas across Boulder, Jefferson, and Adams Counties, affecting cities such as Louisville, Superior, Broomfield, Lafayette, Arvada, and Westminster. A wet spring followed by a hot, dry summer and fall created dry fuel conditions that, combined with strong winds, accelerated the fire’s spread. As Colorado’s most destructive wildfire, it burned over 6,200 acres, destroyed 1,084 homes, damaged 149 others, and caused two fatalities. In Boulder County alone, residential damages exceeded \$513 million. Over 30,000 residents were evacuated on the day of the fire. Figure 5 shows the affected evacuation and pre-evacuation zones Forrister et al. (2024a).
@@ Appendix F Selected Questions for Different Perceptions
It is also worth noting that while some variables listed in Table 12 and Table 13 (e.g., observing flames) may appear difficult to obtain directly during a wildfire, they can often be approximated using external sensor data such as UAV imagery or real-time fire spread models. Similarly, intermediate factors influencing decision-making can be inferred from anticipatory behavioral signals, including social media activity or traffic patterns, thereby extending FLARE’s ability to predict individual behavior even before direct wildfire exposure.
@@ 10. How long had you lived at that residence? (Select only one)
Appendix G Comparison with Actual Reasoning Process of Residents
To assess the reasoning processes generated by LLMs, we compare them with open-ended responses provided by survey participants (answered by 206 out of 604 interviewees), which serve as partial ground truth for reasoning processes.
For example, the self-reported (ground truth) reasoning pattern of one resident is:
“I did not believe that I was in an area at risk of wildfires, I did not think that there would be a fire, I observed others staying behind (in person OR on TV, radio, or social media), I wanted or needed to protect (or defend) my property, I wanted or needed to take care of pets or livestock (or other farm animals), I thought that it was safer to stay in place.”
And the reasoning process generated by the LLM for this resident is:
Let me create a logical chain of reasoning to determine if this resident evacuated:
Starting with pre-fire indicators: the resident had high theoretical risk awareness (5/5 rating), was well-prepared with emergency plans, and had lived in the area for over 10 years. These factors might suggest evacuation likelihood.
However, examining actual emergency response: the resident showed no reaction to official notifications, ignored mandatory evacuation orders, and explicitly planned to stay at home. They offered several reasons for staying: they did not believe the area was at risk, wanted to protect their property, and needed to care for pets or livestock.
Comparing to previous examples: Example 1 had moderate risk perception and evacuated; Example 2 had low risk perception but still evacuated. In contrast, this resident exhibited active resistance to evacuation.
Additional supporting evidence includes a large household (5 adults), multiple specific reasons for staying, and a consistent pattern of ignoring warnings.
Final analysis: the resident’s behavior demonstrates a clear and deliberate choice to stay, supported by multiple justifications and consistent dismissal of evacuation orders.
Conclusion: NO — this resident did not evacuate during the wildfire.
We observe that the LLM’s reasoning converges on key motivational themes that align with the resident’s own explanation. This alignment demonstrates the framework’s ability to capture nuanced behavioral logic, indicating that the generated reasoning is not only predictive but also interpretable and psychologically grounded.